The base of an isosceles triangle is 12in, less than the length of the equal sides. If the perimeter of the triangle is 39in., find the length of each of the sides.
Answer 17 for each side and 5 for the base. Please explain.

Answer by rapaljer(4671)

If the triangle is isosceles, this means that TWO sides are equal. If the base is 12, then let the two equal sides be x.

The base is x-12

The sum of the three sides is
x+x+x-12 = 39
3x-12=39
3x =39+12
3x=51
x=51/3
x=17 inches for each equal side
x-12= 5 inches = base

R^2